Flamo:

Find the distance between the points (2.7, 5.1) and (3, 4.7).

Flamo:

Same thing but with Decimals?

Flamo:

O, Its the Distance

dude:

Yes Distance formula =\(\sqrt{(x_2 -x_1)^2 + (y_2- y_1)^2}\)

Flamo:

o

Flamo:

Ye that's the formula.

dude:

Yeah so just substitute the points into the equation

Flamo:

What Points go first? I always get Confused on that...

Flamo:

4.7 and 5.1?

Flamo:

Then 3 and 2.7?

Shadow:

You can always start by labeling them on a piece of paper. First label the x points, then the y points. Then label the second set of points as 2, and then the first set as 1.

Flamo:

(2.7, 5.1) and (3, 4.7). X Y X Y

Flamo:

Right?

Hero:

(2.7, 5.1) and (3, 4.7). Xₗ Yₗ X₂ Y₂

Flamo:

o

Flamo:

\[d = \sqrt{(3 - 2.7)^2 + (3 - 4.7)^2}\]

Flamo:

Right?

Hero:

Where'd you get two 3's from?

Flamo:

...

Flamo:

Woops.

Flamo:

\[d = \sqrt{(3-2.7)^2 + (4.7-5.1)^2}\]

Flamo:

There?

Flamo:

\[d = \sqrt{(0.3)^2 + (-0.4)^2}\]

Flamo:

\[d = \sqrt{0.09 + (-0.16)}\]

Flamo:

\[d = \sqrt{-0.7}\]

Flamo:

... That's incorrect right @Hero?

Hero:

It's not correct. Any time you square a number, you should have positive numbers afterwards.

Flamo:

o